OpenStudy (imamermaid29):

Which ordered pair is the solution to the system of equations? x + y = 4 x - y = -6 A. (2, 2) B. (1, −7) C. (−1, 5) D. (−2, 4)

OpenStudy (imamermaid29):

@green_1 do u know this by any chance ?

OpenStudy (green_1):

i think it is c because -1+5=4 and -1-5=-6

OpenStudy (imamermaid29):

okay thanks :) @green_1

OpenStudy (green_1):

need anymore help?

OpenStudy (imamermaid29):

yes @green_1

OpenStudy (imamermaid29):

Which ordered pair is the solution to the system of equations? 2x - y = -5 x + 3y = 22 A. (1, 7) B. (4, 9) C. (7, 19) D. (8, 21)

OpenStudy (green_1):

it would be A

OpenStudy (imamermaid29):

i think i have like 2-3 more could you help? @green_1

OpenStudy (green_1):

sure

OpenStudy (imamermaid29):

Which ordered pair is the solution to the system of equations? x+3y = 12 -4x + 3y = -3 A. (0, 4) B. (6, 2) C. (3, 3) D. (0, −1)

OpenStudy (anonymous):

C. seems to make the most sense to me to me cause it fits the 1st equation and i can clearly tell that the others don't. so yeah it's C.

OpenStudy (imamermaid29):

Thanks :) @raventayler13

OpenStudy (anonymous):

welcome

OpenStudy (imamermaid29):

Tickets to the zoo cost $14 for adults and $10 for children. A group of 12 people went to the zoo, and the tickets cost $140. The system of equations models this situation, where x is the number of adults and y is the number of children. x+y = 12 14x + 10y = 140 How many adults and how many children were in the group? A. 5 adults and 7 children B. 6 adults and 6 children C. 7 adults and 5 children D. 8 adults and 4 children

OpenStudy (green_1):

i think it is a

OpenStudy (anonymous):

yes it is A.

OpenStudy (imamermaid29):

thanks @green_1 & @raventayler13

OpenStudy (anonymous):

welcome @ImaMermaid29

OpenStudy (green_1):

just tag me and i will try to help if im online

OpenStudy (imamermaid29):

A container holds a combination of nickels and quarters. There are 156 coins in all, and their total value is $21.60. The system of equations models this situation, where x is the number of nickels and y is the number of quarters. x + y = 156 0.05x + 0.25y = 21.6 How many nickels and quarters are in the container? A. 118 nickels and 38 quarters B. 106 nickels and 50 quarters C. 87 nickels and 69 quarters D. 72 nickels and 84 quarters

OpenStudy (imamermaid29):

thats my last question @green_1 & @raventayler13

OpenStudy (anonymous):

it's C.

OpenStudy (imamermaid29):

thank you so so much!! :) @raventayler13

OpenStudy (anonymous):

Dōitashimapellete

OpenStudy (anonymous):

lol auto correct it is Dōitashimapellete

OpenStudy (anonymous):

damn thing that's not it it's supposed to be saying your welcome in japanese but it keeps correcting it oh well i guess you-if you want to-will have to look it up on google translate. lol XD

OpenStudy (imamermaid29):

Dōitashimapellete @raventayler13

OpenStudy (imamermaid29):

damn it it did auto correct XD

OpenStudy (anonymous):

Dōitashima is what it got correct but pellete was corrected. lol

OpenStudy (anonymous):

see did it again!

OpenStudy (imamermaid29):

ik XD

OpenStudy (anonymous):

and I can only say 3 things in japanese. so i need to work on it but i guess if i watch enough corpse party i might get better.

OpenStudy (imamermaid29):

lol

OpenStudy (anonymous):

wanna know what I can say?

OpenStudy (imamermaid29):

sure :)

OpenStudy (anonymous):

ok I can say thank you, your welcome and, i can also say don't look. I learned the last one just recently.

OpenStudy (imamermaid29):

looks like i need to step up my game

OpenStudy (anonymous):

really? don't you know any japanese?

OpenStudy (imamermaid29):

i know hello & what are those XD

OpenStudy (anonymous):

well then lol XD

OpenStudy (imamermaid29):

yeah ik ... hey just message me okay:) @raventayler13

OpenStudy (anonymous):

ok ... wait right now?

OpenStudy (imamermaid29):

yeah @raventayler13 or when ever

OpenStudy (anonymous):

ok lol XD